# Understanding Domain-Size Generalization in Markov Logic Networks

### v1

@@ 7.2 Methodology
We compare the generalization behavior of standard generative weight learning to different methods that also reduce parameter variance: L1 regularization, L2 regularization, and DA-MLNs. The structures of the MLNs we use are adopted from Van Haaren et al. [27], who introduced a Lifted Structure Learning (LSL) approach. LSL ensures that the learned structures are liftable and learnable in practice. The Nations dataset, with over 160 predicates, presents an infeasibly large search space of possible clauses for LSL. Hence, we use a hand-crafted MLN of 50 formulas. For weight learning, we employ Lifted Generative Learning [27]. This allows us to compare the exact target-dataset likelihoods, which is the natural evaluation measure for generative learning [4, 12, 17, 27] and is also best suited for validating our theoretical results.
To provide reliable results we generate 20 training samples for weight learning and 5 target samples for each of the sizes we want to generalize to. For generating a training sample, we uniformly sample a subset from a specific type $\tau$ of constant: We sample 20 persons for FS, 50 persons for IMDB and WebKB, and 5 nations for the Nations dataset. Now, let $\mathrm{I}$ denote the set of the sampled constants. In the training sample, we then include all the ground atoms $R(a_{1},...a_{k})$ where all the domain constants of type $\tau$ in $\{a_{1},...,a_{k}\}$ are included in $\mathrm{I}$. The process for generating target samples follows a similar approach.
For standard generative weight learning and DA-MLNs, we then learn the weights on each training sample and compute the log-likelihood of each target dataset. For L1 and L2 regularization, to find the best regularization parameter $\lambda$, we perform hyperparameter tuning on the values between $10^{-2}$ and $10^{2}$ on the smallest target samples  [^4] Hence, the results on the smallest target samples are slightly biased for L1 and L2. However, for larger target sizes, no such bias exists.. As our metric to compare the different approaches, we measure how much the target set log-likelihood improves in comparison to no regularization. This metric measures how well our ML estimate is w.r.t. equation (3), i.e., the ML estimate that takes into account the fact that the observed data came from a larger relational structure.
@@ 7.3 Results
Figure 1 shows the difference between the average log-likelihood obtained with the regularization approaches and the one obtained without regularization. For each of the four datasets, methods that reduce parameter variance consistently improve target set likelihood by several orders of magnitude (except for L2 regularization on FS). This effect is more pronounced as the target set size grows. Among the methods that reduce parameter variance, L1 and L2 regularization have similar performances. DA-MLNs outperform L1 and L2 on the FS dataset, but underperform on Nations and IMDB, while producing similar results on WebKB. Note that MLNs learned with L1 and L2 regularization are not domain-aware and work with the same parameters across domain sizes. While our theoretical results provide bounds depending on the target domain size, it is unclear whether reducing the variance based on the target set size is necessary in practice. This can be observed in the relative under-performance of DA-MLNs on some of the datasets.
@@ 9 Acknowledgments
Proof (of Proposition 4)
@@
Proof (of Theorem 6.1)

### v3

@@ 7.2 Methodology
We compare the generalization behavior of standard generative weight learning to different methods that also reduce parameter variance: L1 regularization, L2 regularization, and DA-MLNs. The structures of the MLNs we use are adopted from Van Haaren et al. [27], who introduced a Lifted Structure Learning (LSL) approach. LSL ensures that the learned structures are liftable and learnable in practice. The Nations dataset, with over 160 predicates, presents an infeasibly large search space of possible clauses for LSL. Hence, we use a hand-crafted MLN of 50 formulas. For weight learning, we employ Lifted Generative Learning [27]. This allows us to compare the exact target set likelihoods, which is the natural evaluation measure for generative learning [4, 13, 18, 27] and is also best suited for validating our theoretical results.
To provide reliable results we generate 20 training set for weight learning and 5 target sets for each of the sizes we want to generalize to. For generating a training set, we uniformly sample a subset from a specific type $\tau$ of constant: We sample 20 persons for FS, 50 persons for IMDB and WebKB, and 5 nations for the Nations dataset. Now, let $\mathrm{I}$ denote the set of the sampled constants. In the training set, we then include all the ground atoms $R(a_{1},...a_{k})$ where all the domain constants of type $\tau$ in $\{a_{1},...,a_{k}\}$ are included in $\mathrm{I}$. The process for generating target sets follows a similar approach.
For standard generative weight learning and DA-MLNs, we then learn the weights on each training set and compute the log-likelihood of each target set. For L1 and L2 regularization, to find the best regularization parameter $\lambda$, we perform hyperparameter tuning on the values between $10^{-2}$ and $10^{2}$ on the smallest target sets [^5] Hence, the results on the smallest target sets are slightly biased for L1 and L2. However, for larger target set sizes, no such bias exists.. As our metric to compare the different approaches, we measure how much the target set log-likelihood improves in comparison to no regularization. This metric measures how well our ML estimate is w.r.t. equation (3), i.e., the ML estimate that takes into account the fact that the observed data came from a larger relational structure.
@@ 7.3 Results
Figure 1 shows the difference between the average log-likelihood obtained with the regularization approaches and the one obtained without regularization. For each of the four datasets, methods that reduce parameter variance consistently improve target set likelihood by several orders of magnitude (except for L2 regularization on FS). This effect is more pronounced as the target set size grows. Among the methods that reduce parameter variance, L1 and L2 regularization have similar performances. DA-MLNs outperform L1 and L2 on the FS dataset, but underperform on Nations and IMDB, while producing similar results on WebKB. Note that MLNs learned with L1 and L2 regularization are not domain-aware and work with the same parameters across domain sizes. Thus, it is unclear whether domain-aware parameter variance reduction methods are generally preferable to domain-unaware methods in practice. This can be observed in the relative under-performance of DA-MLNs on some of the datasets.
@@ 9 Acknowledgments
Proof (the lower bound proof for Proposition 3)
@@
Proof (the lower bound proof for Theorem 6.1)